### v3

# Uncertainty Makes It Stable: Curiosity-Driven Quantized Mixture-of-Experts

Abstract
Deploying neural networks on edge devices requires preserving accuracy under aggressive quantization while ensuring stable, predictable inference. We introduce a curiosity-driven quantized Mixture-of-Experts (MoE) framework that combines heterogeneous experts (BitNet ternary, 1–16 bit BitLinear, and post-training quantized models) with Bayesian epistemic uncertainty-based routing. Evaluated on three audio classification benchmarks (ESC-50, Quinn, UrbanSound8K), 4-bit quantization preserves 99.9% of full-precision F1 (0.858 vs 0.859) with 4 compression and 31% energy savings over 8-bit, while both achieve statistical parity with full precision ().
Curiosity-driven routing simultaneously improves accuracy and stability: on Quinn, F1 increases from 0.802 to 0.809 while cross-fold variance drops by 85% (, Levene’s test), with reductions of 50–94% across datasets. The routing is self-organizing: the high-precision 8-bit expert automatically receives the most uncertain samples (20% lower confidence, ), while lightweight experts handle easier inputs. Datasets with already-low baseline variance show no artificial stability gain, confirming the mechanism targets genuine epistemic uncertainty rather than overfitting routing decisions. At 1.2M parameters, the framework provides interpretable, precision-aware routing suitable for safety-sensitive edge deployments where both accuracy and predictability are critical.
1 Introduction
Deploying deep neural networks on resource-constrained edge devices requires balancing model accuracy, computational efficiency, and energy consumption. While modern architectures achieve impressive performance on high-end hardware, their memory footprint and inference costs often prohibit deployment on mobile and embedded systems where real-time audio processing is most valuable Jacob et al. (2018). This challenge is particularly acute for environmental sound classification, where models must operate continuously on battery-powered devices Torija et al. (2014).
Quantization has emerged as a leading compression approach, reducing numerical precision from 32-bit floating-point to lower bit-widths Jacob et al. (2018). However, aggressive quantization frequently degrades accuracy, especially when applying uniform bit-widths across all network components. Recent work on Mixture of Experts (MoE) architectures suggests heterogeneous model ensembles can improve performance Souli and Lachiri (2018), yet existing MoE frameworks rely on fixed routing policies that fail to adapt to input complexity and quantization-induced uncertainty.
We introduce a curiosity-driven quantized Mixture of Experts framework addressing these limitations through three key innovations. First, we implement heterogeneous quantization by deploying diverse expert types within a single MoE architecture, including BitNet ternary quantization, BitLinear schemes spanning 1 to 16 bits, and post-training quantization (PTQ) with bitwise operations. Second, we propose Bayesian curiosity-driven routing that selects experts based on both prediction confidence and epistemic uncertainty, encouraging exploration of quantization strategies maximizing information gain (Figure 1). Third, we conduct comprehensive efficiency analysis across energy consumption, carbon emissions, and latency stability, metrics critical for sustainable edge deployment.
Our evaluation on ESC-50 Piczak (2015), Quinn Quinn et al. (2022), and UrbanSound8K Salamon et al. (2014) establishes 4-bit as the optimal operating point: 99.9% of 16-bit accuracy (0.858 vs 0.859 F1) with 4 compression and 41% energy savings versus 8-bit. Crucially, curiosity-driven routing reduces MoE F1 variance by 50–94% on challenging datasets (, Levene’s test), with the 8-bit expert automatically receiving the most uncertain samples (20% lower confidence, ). Energy analysis reveals counter-intuitive patterns: 16-bit consumes least energy (0.018 mJ) despite highest precision, while 8-bit requires 2.7 more (0.048 mJ) due to dequantization overhead on Apple M3 CPU. Deployment-phase emissions dominate training by 10,000 for models serving more than 1M inferences.
Contributions: (1) Comprehensive quantization framework evaluating BitNet, BitLinear, and PTQ schemes across 1 to 16 bits for audio classification, establishing 4-bit as optimal accuracy-efficiency operating point. (2) Curiosity-driven routing via Bayesian epistemic uncertainty, achieving 50–94% F1 variance reduction () with automatic adaptive precision allocation. (3) Holistic efficiency assessment including energy consumption, carbon emissions, and statistical significance testing across three diverse audio benchmarks, demonstrating that simple 4-bit quantized models outperform complex MoE architectures for most edge deployments. (4) Hardware-specific analysis revealing CPU bottlenecks limiting quantization speedups, with deployment recommendations for five constraint scenarios: latency-critical, balanced, maximum accuracy, energy-constrained, and stable real-time systems.
2 Related Work
2.1 Audio Representation Learning
Environmental sound classification traditionally relied on hand-crafted features such as Mel-frequency cepstral coefficients (MFCCs) and Mel spectrograms Davis and Suresh (2018); Stevens et al. (1937). Recent advances have shifted toward learned representations, with pre-trained models like VGGish Hershey et al. (2016), OpenL3 Cramer et al. (2019), and AudioCLIP Guzhov et al. (2021) demonstrating strong transfer learning capabilities Mesaros et al. (2018). We leverage pre-trained visual models (EfficientNet-B3 Tan and Le (2019), MobileNet-v3 Howard et al. (2019)) applied to Mel spectrograms, exploiting structural similarities between time-frequency representations and natural images.
2.2 Neural Network Quantization
Quantization reduces memory footprint and computational cost by representing weights and activations with lower numerical precision. Post-training quantization (PTQ) converts pre-trained FP32 models to INT8 or lower bit-widths without retraining Jacob et al. (2018); Oh et al. (2022), while quantization-aware training (QAT) incorporates quantization during training for more aggressive compression Kaziha and Bonny (2020); Hernández et al. (2024). Recent advancements include layer-wise adaptive quantization Liu and Chen (2024), ternary quantization Ma et al. (2024) using values, and binary quantization (1-bit) with gradient approximation Gao et al. (2024). Mixed-precision quantization assigns heterogeneous bit-widths across layers Kim et al. (2020), while bitwise operations enable ultra-low-power inference Hubara et al. (2016).
However, most research focuses on quantizing full models Hubara et al. (2016), with less attention to quantization’s impact on embedding quality for edge applications Bhardwaj et al. (2019); Esser et al. (2020). Existing approaches typically apply uniform quantization policies to entire networks, leaving unexplored dynamic quantization strategy selection based on input characteristics, a gap our work addresses through heterogeneous MoE architectures with curiosity-driven routing.
2.3 Mixture of Experts Architectures
Mixture of Experts (MoE) architectures improve model capacity by partitioning computation across specialized sub-networks (experts), with a gating network determining expert activation based on input features Pavlitskaya and Harmeling (2022); Sharma and Tripathi (2019). Sparse MoE models activate only a subset of experts per input, reducing inference cost while maintaining large total capacity Zhou et al. (2022); Riquelme et al. (2021). Routing mechanisms have evolved from simple learned gates to sophisticated assignment strategies with load-balancing constraints and differentiable top- routing Puigcerver et al. (2024). However, existing MoE frameworks do not exploit quantization heterogeneity, treating all experts as operating at the same numerical precision. Furthermore, routing decisions typically rely solely on prediction confidence without considering epistemic uncertainty, which could guide exploration when model confidence is unreliable.
2.4 Uncertainty-Driven Learning
Bayesian deep learning quantifies epistemic uncertainty arising from limited training data Zeevi et al. (2025). Monte Carlo dropout approximates Bayesian inference by sampling multiple forward passes Kendall and Gal (2017), providing uncertainty estimates without modifying architecture. Curiosity-driven learning leverages information-theoretic principles to encourage exploration in reinforcement learning Pathak et al. (2017). Our work bridges this gap by introducing Bayesian curiosity-driven routing for quantized MoE architectures, enabling adaptive selection of heterogeneous quantization schemes based on epistemic uncertainty Ordoñez (2025); Celi (2025).
2.5 Quantization for Edge Deployment
Recent work has demonstrated successful deployment of quantized models on edge devices, but several gaps remain. Most studies report accuracy metrics without comprehensive energy consumption or carbon emission measurements Bhardwaj et al. (2019). Hardware-specific quantization effects are often overlooked; CPU architectures lacking native INT8 acceleration may exhibit counter-intuitive energy patterns where lower bit-widths consume more power due to dequantization overhead. Additionally, latency variance is rarely analyzed despite its importance for battery-constrained devices.
Our work addresses these gaps by providing comprehensive efficiency analysis across energy, carbon emissions, and statistical significance testing. We demonstrate that 4-bit quantization achieves optimal accuracy-efficiency trade-offs, and that curiosity-driven routing reduces F1 variance by 50–94% on challenging datasets () with adaptive precision allocation, validated through rigorous statistical testing with Bonferroni correction.
3 Methodology
Our methodology comprises three stages: audio representation extraction, heterogeneous quantization, and curiosity-driven expert routing. We generate Mel spectrogram representations from raw audio and extract embeddings using pre-trained convolutional networks. These embeddings are processed through multiple quantized expert classifiers employing diverse compression schemes. A Bayesian router dynamically selects experts based on prediction confidence and epistemic uncertainty.
3.1 Audio Representation and Embedding Extraction
We represent audio signals as Mel spectrograms, transforming raw waveforms into time-frequency representations aligned with human auditory perception. Given an audio signal , we compute the Short-Time Fourier Transform (STFT) with window size 2048 and hop length 512, generating magnitude spectrogram . We apply triangular Mel-scale filters to map linear frequencies to perceptual frequencies:
We extract 1024-dimensional embeddings from the penultimate layer of pre-trained EfficientNet-B3 and MobileNet-v3 models. These visual feature extractors, trained on ImageNet, generalize effectively to spectrogram analysis due to structural similarities between time-frequency representations and natural images Hershey et al. (2016). The embeddings encode high-level acoustic patterns providing semantic audio content representation.
3.2 Heterogeneous Quantization Framework
We implement three distinct quantization schemes to create diverse experts:
BitLinear Quantization. We apply uniform -bit quantization to weight matrices using symmetric quantization with learnable scale parameters. For weight tensor :
where is the symmetric quantization scale. We evaluate BitLinear experts with bits, denoted as Q-Base models.
BitNet Ternary Quantization. BitNet restricts weights to , enabling multiplication-free inference:
where is a learned threshold. This reduces model size by 16 compared to INT8 while enabling efficient bitwise operations.
Post-Training Quantization (PTQ) with Bitwise Operations. Our PTQ scheme (Q8-Base-PTQ) leverages bitwise popcount operations for ultra-efficient inference. We quantize activations and weights to binary representations and compute layer outputs:
where and are binarized inputs and weights, denotes XOR, and is feature dimension. This replaces expensive multiply-accumulate operations with single-cycle bitwise instructions. We use fixed threshold for binarization.
3.3 Curiosity-Driven Mixture of Experts
Our MoE architecture consists of expert networks employing different quantization schemes. A gating network computes routing probabilities based on input embeddings :
where is temperature controlling routing sharpness. We select top- experts and compute final output as weighted combination:
To incorporate epistemic uncertainty, we augment the router with a Bayesian curiosity term. We compute predictive entropy across experts:
where is average predicted probability for class . High entropy indicates disagreement among quantization strategies, suggesting exploration may be beneficial. We modify routing probabilities by adding a curiosity bonus:
where controls curiosity-driven exploration strength. This encourages selecting experts with diverse predictions when model uncertainty is high.
Hyperparameter Sensitivity. The curiosity strength governs the degree of routing polarization. Too low values () cause router collapse to a single expert (100% Q4), providing no benefit over fixed quantization. Effective values lie in : at , all three experts remain active (42% BitNet, 39% Q4, 19% Q8) with 94% variance reduction; at , routing polarizes to two experts (77% Q4, 23% Q8) achieving best F1 (0.7820.015) with 80% variance reduction.
Theoretical Justification for Stability. We model inference latency as , where is the compute cost determined by expert selection and captures hardware noise (interrupt handling, CPU scheduling). By the law of total variance:
since hardware noise is negligible compared to routing variance. Uniform routing causes frequent switching between experts with different computational costs (e.g., Q4 vs Q8), producing high . Curiosity routing (Eq. 8) amplifies KL divergence, pushing toward one-hot assignments. When a single expert dominates (), switching between cost modes vanishes, directly reducing .
4 Experiments
4.1 Experimental Setup
We evaluate on ESC-50 environmental sound classification (2,000 clips, 50 classes, 5 predefined folds) using 5-fold cross-validation. Each fold trains on 1,600 samples and validates on 400 held-out samples. Performance metrics are reported as mean standard deviation across 5 folds. Statistical significance is assessed using paired t-tests with Bonferroni correction (). Additional validation on Quinn and UrbanSound8K datasets confirms cross-dataset generalization.
4.2 Model Architecture
Our architecture extracts 1024-dimensional embeddings from pre-trained image classifiers (EfficientNet-B3, MobileNet-v3), then processes through an MLP with hidden dimensions [256, 128, 64], ReLU activations, and dropout (). Quantized pipelines incorporate linear dequantization blocks at network input, maintaining quantization parameters throughout forward passes.
4.3 Implementation Details
Router: Three-layer network: Linear(1024, 128), ReLU, Dropout(0.2), Linear(128, 64), ReLU, Linear(64, num_experts) with Xavier initialization. Bayesian routing uses 10 Monte Carlo dropout samples for epistemic uncertainty estimation.
Hyperparameters: Individual models use AdamW (lr=5.79e-4, weight decay=5.13e-3, batch size=64) with hidden dimensions [640, 320] and dropout 0.195. MoE models use Adam (lr=1e-3, weight decay=1e-4, batch size=256). Early stopping patience: 19 epochs (individual), 30 epochs (MoE). Load balancing , temperature . Curiosity routing uses KL divergence strategy with strength (see Section 5.3).
Quantization: BitLinear uses symmetric per-layer scales: for weights, for activations. BitNet learns per-channel scales via gradient descent. Q8-Base-PTQ uses fixed threshold .
All models trained with gradient clipping (max norm=1.0) and class-weighted cross-entropy loss on Apple M3 Max CPU (36GB RAM) using PyTorch 2.0.1 with qnnpack engine.
5 Results
5.1 Cross-Dataset Generalization
We evaluate all models across ESC-50 (environmental sounds), Quinn (acoustic scenes), and UrbanSound8K (urban soundscapes) using 5-fold cross-validation. Table 1 presents comprehensive results as mean±standard deviation.
| Model | Params (M) | ESC-50 F1 | Quinn F1 | Urban8K F1 | Avg F1 |
| Q16-Base | 1.2 | 0.819±0.020 | 0.804±0.005 | 0.952±0.006 | 0.859±0.012 |
| Q4-Base | 1.2 | 0.819±0.021 | 0.805±0.009 | 0.950±0.008 | 0.858±0.014 |
| Q8-Base | 1.2 | 0.818±0.012 | 0.808±0.008 | 0.948±0.010 | 0.858±0.010 |
| FP32-Base† | 1.19 | 0.815±0.018 | 0.803±0.006 | 0.946±0.005 | 0.855±0.011 |
| Q8-Base-PTQ‡ | 1.19 | 0.810±0.015 | 0.802±0.009 | 0.950±0.006 | 0.854±0.011 |
| Q1-Base | 1.2 | 0.806±0.018 | 0.803±0.004 | 0.944±0.009 | 0.851±0.012 |
| BitNet-Q4/8/16-QMoE | 4.99 | 0.791±0.016 | 0.812±0.006 | 0.943±0.003 | 0.849±0.010 |
| BitNet-Q4/8-QMoE-C§ | 3.8 | 0.782±0.015 | 0.805±0.007 | 0.941±0.011 | 0.843±0.011 |
| BitNet-Q4/8/16-QMoE-C§ | 4.99 | 0.788±0.013 | 0.803±0.009 | 0.941±0.008 | 0.844±0.010 |
| Q2-Base | 1.2 | 0.783±0.010 | 0.796±0.011 | 0.921±0.010 | 0.833±0.010 |
| BitNet-Q8/16-PTQ-QMoE-C§ | 4.99 | 0.750±0.022 | 0.807±0.006 | 0.940±0.008 | 0.832±0.014 |
| BitNet-Q8/16-PTQ-QMoE | 4.99 | 0.780±0.029 | 0.797±0.019 | 0.937±0.013 | 0.820±0.020 |
| BitNet-Q8PTQ-QMoE-C§ | 2.59 | 0.765±0.029 | 0.795±0.009 | 0.935±0.010 | 0.832±0.018 |
| BitNet-Q8PTQ-QMoE | 2.59 | 0.740±0.017 | 0.793±0.012 | 0.929±0.013 | 0.801±0.013 |
| BitNet-Base | 1.2 | 0.705±0.013 | 0.768±0.007 | 0.906±0.005 | 0.793±0.009 |
Q16-Base achieves the best average F1-score (0.859±0.012), matching FP32-Base† (0.855±0.011). Notably, 4-bit and 8-bit quantization maintain 99.9% of 16-bit accuracy (0.858 F1) with only 1.2M parameters, providing substantial computational savings. Performance varies by dataset: UrbanSound8K shows highest accuracy (0.906-0.952 F1) due to structured urban sounds, while ESC-50 (0.705-0.819) and Quinn (0.768-0.812) prove more challenging.
MoE architectures show mixed generalization. BitNet-Q4/8/16-QMoE achieves best Quinn performance (0.812±0.006) but underperforms simple quantized models on ESC-50 and UrbanSound8K, suggesting expert routing benefits dataset-specific adaptation but introduces overhead for simpler tasks. Low standard deviations (±0.004-0.033) indicate stable performance, with quantization-aware training (Q4/Q8/Q16-Base) consistently matching or exceeding post-training quantization (Q8-Base-PTQ‡).
5.2 Quantization Bit-Width Ablation
We systematically evaluate quantization precision across 1, 2, 4, 8, and 16-bit schemes. Table 2 presents performance-efficiency trade-offs on ESC-50, Quinn, and UrbanSound8K, with all models maintaining identical 1.2M-parameter architecture to isolate bit-width effects.
| Bits | Dataset | Params | F1-Score | % of | Latency | Energy | RAM |
| (M) | 16-bit | (ms) | (mJ) | (GB) | |||
| 1 | ESC-50 | 1.2 | 0.806±0.018 | 98.3 | 1103.55±49.50 | 0.054±0.031 | 0.85±0.03 |
| 2 | 1.2 | 0.783±0.010 | 95.5 | 1117.79±46.82 | 0.034±0.018 | 0.84±0.06 | |
| 4 | 1.2 | 0.819±0.021 | 99.9 | 1126.03±51.66 | 0.034±0.023 | 0.86±0.04 | |
| 8 | 1.2 | 0.818±0.012 | 99.8 | 1098.21±19.35 | 0.048±0.021 | 0.93±0.01 | |
| 16 | 1.2 | 0.819±0.020 | 100.0 | 1459.19±10.58 | 0.018±0.003 | 0.95±0.00 | |
| 1 | Quinn | 1.2 | 0.803±0.004 | 99.8 | 1197.66±14.50 | 0.122±0.021 | 1.34±0.11 |
| 2 | 1.2 | 0.796±0.011 | 99.0 | 1195.76±34.13 | 0.134±0.040 | 1.33±0.17 | |
| 4 | 1.2 | 0.805±0.009 | 100.1 | 1199.26±36.43 | 0.167±0.045 | 1.26±0.30 | |
| 8 | 1.2 | 0.808±0.008 | 100.4 | 1209.00±6.67 | 0.170±0.040 | 1.09±0.10 | |
| 16 | 1.2 | 0.804±0.005 | 100.0 | 1129.87±41.36 | 0.065±0.005 | 1.32±0.17 | |
| 1 | UrbanSound8K | 1.2 | 0.944±0.009 | 99.2 | 1187.49±9.24 | 0.128±0.025 | 1.30±0.14 |
| 2 | 1.2 | 0.921±0.010 | 96.7 | 1210.45±11.38 | 0.197±0.051 | 1.31±0.21 | |
| 4 | 1.2 | 0.950±0.008 | 99.7 | 1235.90±17.80 | 0.213±0.033 | 1.10±0.40 | |
| 8 | 1.2 | 0.948±0.010 | 99.5 | 1203.95±2.90 | 0.217±0.066 | 1.03±0.23 | |
| 16 | 1.2 | 0.952±0.006 | 100.0 | 1178.25±8.19 | 0.126±0.038 | 1.06±0.30 |
Performance vs. Bit-Width Trade-offs. Our ablation reveals 4-bit and 8-bit as optimal operating points. 4-bit achieves 99.9% of 16-bit performance (0.819 vs 0.819 F1 on ESC-50) while reducing model size by 4. The performance cliff occurs at 2-bit (95.5%, 0.783 F1), reflecting insufficient precision for weight distributions. Surprisingly, 1-bit maintains 98.3% performance (0.806 F1), suggesting extreme compression preserves discriminative information with quantization-aware training.
Energy-Efficiency Analysis. Energy consumption exhibits a U-shaped relationship with bit-width. 16-bit consumes least energy (0.018 mJ on ESC-50), followed by 4-bit (0.034 mJ). Counter-intuitively, 8-bit requires 2.7 more energy (0.048 mJ) despite fewer bits, likely due to dequantization overhead and non-optimized INT8 operations on Apple M3 CPU. The optimal point balances accuracy and efficiency: 4-bit provides 99.9% accuracy with 31% better energy efficiency than 8-bit.
Latency and Dataset Sensitivity. Inference latency shows hardware-specific patterns: 8-bit achieves lowest latency on ESC-50 (1098±19 ms), suggesting better CPU optimization for byte-aligned operations. UrbanSound8K demonstrates exceptional quantization robustness, maintaining 99% performance down to 1-bit due to structured acoustic categories. ESC-50’s diverse environmental sounds show greater sensitivity (2-bit: 95.5%), suggesting quantization tolerance correlates with inter-class separability.
5.3 Mixture-of-Experts with Curiosity-Driven Routing
We compare heterogeneous quantized MoE architectures using uniform top-k routing against Bayesian epistemic uncertainty-based selection. Table 3 presents performance-efficiency trade-offs across seven configurations, aggregated over ESC-50, Quinn, and UrbanSound8K.
| Model | Type | Params | F1 | Latency | Energy | CO2 | Eff |
| (M) | (ms) | (mJ) | (g) | ||||
| BitNet-Q4/8/16-QMoE | MoE | 4.99 | 0.849±0.010 | 1335±230 | 0.621±0.227 | 228±84 | 1.37 |
| BitNet-Q8/16-PTQ-QMoE | MoE | 4.99 | 0.820±0.025 | 1219±69 | 0.222±0.096 | 82±36 | 3.70 |
| BitNet-Q8PTQ-QMoE | MoE | 2.59 | 0.801±0.013 | 1172±46 | 0.207±0.101 | 76±37 | 3.87 |
| BitNet-Q4/8-QMoE-C§ | MoE-C | 3.80 | 0.843±0.011 | 1278±29 | 0.558±0.145 | 205±53 | 1.51 |
| BitNet-Q4/8/16-QMoE-C§ | MoE-C | 4.99 | 0.844±0.010 | 1281±35 | 0.529±0.144 | 195±53 | 1.59 |
| BitNet-Q8/16-PTQ-QMoE-C§ | MoE-C | 4.99 | 0.832±0.014 | 1251±39 | 0.281±0.093 | 103±34 | 2.96 |
| BitNet-Q8PTQ-QMoE-C§ | MoE-C | 2.59 | 0.832±0.018 | 1225±38 | 0.260±0.108 | 96±40 | 3.20 |
Performance-Efficiency Trade-offs. BitNet-Q4/8/16-QMoE achieves highest F1-score (0.849±0.010) but consumes 3 more energy (0.621 mJ) than the most efficient configuration. BitNet-Q8PTQ-QMoE achieves best efficiency (3.87 F1/mJ) with only 0.207 mJ while maintaining competitive 0.801 F1, a 5.6% accuracy drop for 67% energy savings.
Curiosity-Driven vs. Uniform Routing. Curiosity routing provides dataset-dependent stability: 80% variance reduction on ESC-50 (), 85% on Quinn (), but no benefit on UrbanSound8K (baseline std = 0.007). See Table 4 for per-dataset corrected results. Curiosity routing’s primary value is adaptive precision allocation and prediction stability rather than raw accuracy improvement.
Stability. Curiosity-driven routing collapses routing distributions toward deterministic expert assignments, eliminating cost-mode switching. This reduces F1 variance across folds by 50–94% on ESC-50/Quinn (, Levene’s test) and latency variance (23029 ms std), critical for real-time audio processing requiring predictable response times.
Environmental Impact. CO2 emissions correlate strongly with energy (r=0.97). BitNet-Q4/8/16-QMoE emits 228±84 g per inference versus BitNet-Q8PTQ-QMoE’s 76±37 g, a 3 reduction. Curiosity routing adds 16% emission overhead (150 vs 129 g) due to Monte Carlo sampling. At scale (1M daily inferences), deploying BitNet-Q8PTQ-QMoE over BitNet-Q4/8/16-QMoE reduces annual CO2 by approximately 55 kg.
Deployment Recommendations. (1) Maximum Accuracy: BitNet-Q4/8/16-QMoE (0.849 F1) for abundant resources. (2) Balanced: BitNet-Q4/8-QMoE-C (0.843 F1, 0.558 mJ) for general use. (3) Energy-Constrained: BitNet-Q8PTQ-QMoE (0.801 F1, 0.207 mJ) for edge devices. (4) Real-Time: BitNet-Q8/16-PTQ-QMoE-C (0.832 F1, 38 ms std) for timing guarantees.
Comparison to Single Models. MoE architectures show 1-7% lower accuracy than top single models (Q16-Base: 0.859 F1, Q8-Base: 0.858 F1) but offer advantages in interpretability and extensibility. While simple 4–8 bit quantization of monolithic models remains preferable for raw accuracy, MoE-Curiosity provides per-sample precision adaptation and interpretable routing critical for safety-sensitive applications (e.g., medical diagnosis, autonomous vehicles) where knowing why a particular expert was selected matters.
Per-Dataset Curiosity Results. Table 4 presents per-dataset results across two settings. ESC-50 and Quinn show 50–94% variance reduction, while UrbanSound8K’s already-low baseline variance (std=0.007) leaves no room for improvement, confirming curiosity routing targets genuine epistemic uncertainty. On Quinn, F1 improves from 0.802 to 0.809 while variance drops by 85% at . At on ESC-50, curiosity routing achieves F1=0.7820.015 with 80% variance reduction versus baseline MoE (0.7770.032).
| Dataset | Configuration | F1 | Std | Var. Red. |
| ESC-50 | Baseline MoE (uniform) | 0.777 | 0.032 | — |
| Curiosity | 0.754 | 0.008 | 94% | |
| Curiosity | 0.782 | 0.015 | 80% | |
| Quinn | Baseline MoE (uniform) | 0.802 | 0.011 | — |
| Curiosity | 0.809 | 0.004 | 85% | |
| Curiosity | 0.805 | 0.007 | 50% | |
| Urban8K | Baseline MoE (uniform) | 0.938 | 0.007 | — |
| Curiosity | 0.940 | 0.011 | 132% | |
| Curiosity | 0.941 | 0.011 | 144% |
Confidence-Based Expert Specialization. Analysis of routing decisions (, N=2000 samples) reveals that curiosity routing automatically allocates precision based on sample difficulty (Figure 2). The Q8 (8-bit) expert receives 386 samples (19.3%) with average confidence 0.577, while Q4 receives 777 samples (38.9%, confidence 0.725) and BitNet receives 837 samples (41.9%, confidence 0.719). The Q8 expert receives samples with 20.0% lower confidence than BitNet/Q4 combined (0.577 vs 0.722, , ). This demonstrates that Eq. 8 drives adaptive precision allocation without manual engineering: uncertain inputs are automatically routed to the highest-precision expert where reliability matters most.
Alpha Sensitivity. Curiosity strength controls the routing distribution shape. At , all three experts remain active (42% BitNet, 39% Q4, 19% Q8), achieving best stability (94% variance reduction on ESC-50) at the cost of lower F1 (0.7540.008). At , routing polarizes to two experts (77% Q4, 23% Q8, 0% BitNet), achieving best F1 (0.7820.015) with 80% variance reduction. Quinn confirms the pattern (85%/50% reduction at ). Both settings validate the curiosity mechanism: the router consistently allocates Q8 to harder samples regardless of .
5.4 Inference Latency and Memory Footprint
We benchmark inference latency and peak RAM usage across all configurations. Table 5 presents measurements averaged over ESC-50, Quinn, and UrbanSound8K using 5-fold cross-validation on Apple M3 Max CPU.
| Model | Params | Latency | Speedup | RAM | F1-Score |
| (M) | (ms) | (GB) | |||
| Q1-Base | 1.2 | 1162.90±30.25 | 1.08 | 1.16±0.10 | 0.851 |
| Q8-Base | 1.2 | 1170.39±11.94 | 1.07 | 1.02±0.15 | 0.858 |
| BitNet-Q8PTQ-QMoE | 2.59 | 1171.82±46.48 | 1.07 | 1.01±0.22 | 0.801 |
| Q2-Base | 1.2 | 1174.67±34.09 | 1.07 | 1.16±0.16 | 0.833 |
| BitNet-Base | 1.2 | 1186.44±32.02 | 1.06 | 1.16±0.11 | 0.793 |
| Q4-Base | 1.2 | 1187.06±37.92 | 1.06 | 1.07±0.29 | 0.858 |
| BitNet-Q8/16-PTQ-QMoE | 4.99 | 1218.71±69.24 | 1.03 | 1.03±0.25 | 0.820 |
| BitNet-Q8PTQ-QMoE-C | 2.59 | 1224.91±37.97 | 1.02 | 0.95±0.20 | 0.832 |
| BitNet-Q8/16-PTQ-QMoE-C | 4.99 | 1251.06±38.69 | 1.00 | 0.92±0.17 | 0.832 |
| FP32-Base† | 1.19 | 1254.28±41.57 | 1.00 | 1.17±0.19 | 0.855 |
| Q16-Base | 1.2 | 1255.77±25.10 | 1.00 | 1.11±0.20 | 0.859 |
| Q8-Base-PTQ‡ | 1.19 | 1263.82±32.53 | 0.99 | 1.08±0.18 | 0.854 |
| BitNet-Q4/8-QMoE-C | 3.8 | 1277.65±29.36 | 0.98 | 0.86±0.13 | 0.843 |
| BitNet-Q4/8/16-QMoE-C | 4.99 | 1280.99±34.55 | 0.98 | 0.89±0.14 | 0.844 |
| BitNet-Q4/8/16-QMoE | 4.99 | 1335.48±230.47 | 0.94 | 0.86±0.21 | 0.849 |
Latency-Accuracy Trade-offs. Q1-Base achieves fastest inference (1163±30 ms, 1.08× speedup) despite modest accuracy (0.851 F1), outperforming Q8-Base (1170 ms), Q4-Base (1187 ms), and FP32-Base† (1254 ms). This suggests 1-bit operations benefit from optimized bitwise implementations on Apple M3, avoiding floating-point multiplications. Top F1 models (Q16-Base: 0.859, Q4/Q8-Base: 0.858) exhibit 1170-1256 ms latency, within 7-8% of fastest, indicating framework overhead dominates arithmetic precision for compact 1.2M-parameter models on CPU.
MoE Characteristics and Memory. MoE architectures show 4-17% higher latency than single models, with BitNet-Q4/8/16-QMoE exhibiting 72 larger variance (230 ms std) versus Q8-Base (12 ms std) due to data-dependent routing. Curiosity variants reduce variance to 29-39 ms through epistemic uncertainty-based load balancing. RAM varies 0.86-1.17 GB despite 4 parameter differences (1.2-5.0M), confirming peak RAM is dominated by framework overhead rather than weights; even 1GB RAM devices can accommodate our models.
Latency Stability. Curiosity routing reduces F1 variance by 50–94% on ESC-50/Quinn (, Levene’s test) and latency variance from 230 ms to 29 ms std. The Bayesian router adds 12% overhead (14.6 ms of 123 ms per inference) due to 10 Monte Carlo dropout forward passes, acceptable given 50–94% variance reduction on ESC-50/Quinn ().
Deployment Recommendations. (1) Latency-Critical (1200ms): Q1-Base (1163 ms, 0.851 F1) or Q8-Base (1170 ms, 0.858 F1). (2) Balanced: Q4-Base (1187 ms, 0.858 F1). (3) Maximum Accuracy: Q16-Base (1256 ms, 0.859 F1). (4) Memory-Constrained (900MB): BitNet-Q4/8-QMoE-C (1278 ms, 0.86 GB, 0.843 F1). (5) Stable Real-Time: Q8-Base (1170±12 ms).
5.5 Statistical Significance Testing
We validate findings using paired t-tests on 5-fold cross-validation scores with Bonferroni correction (, = number of comparisons).
Cross-Dataset Generalization. Limited significant differences exist between quantized models and FP32 baseline. Q16-Base significantly outperforms FP32-Base on UrbanSound8K (0.952 vs 0.946 F1, p=0.022) and BitNet-Q8/16-PTQ-QMoE-C on Quinn (0.812 vs 0.803 F1, p=0.021), but no models significantly exceed baseline on ESC-50. This suggests dataset inter-class separability determines whether quantization maintains statistical parity with full precision.
Quantization Bit-Width Effects. Extreme quantization degrades performance significantly: 2-bit on ESC-50 (0.783 vs 0.819 F1, p=0.004) and both 1-bit (p=0.013) and 2-bit (p=0.005) on UrbanSound8K. However, 4-bit and 8-bit maintain statistical parity with 16-bit across all datasets (p0.05), validating 4-bit as optimal and justifying aggressive quantization for deployment.
Curiosity-Driven Routing Benefits. Curiosity routing maintains accuracy equivalence with uniform routing (p0.05 on ESC-50/UrbanSound8K) while delivering substantial stability gains. Curiosity routing reduces F1 variance by 50–94% on ESC-50/Quinn (, Levene’s test). UrbanSound8K shows no stability benefit (baseline std=0.007), confirming the mechanism targets genuine uncertainty. The confidence-based expert specialization is highly significant: Q8 receives samples with 20.0% lower confidence than BitNet/Q4 (, , N=2000), the strongest statistical result in our evaluation.
Latency and Variance. Quantized models show non-significant latency reductions versus FP32 (Q1-Base: 1.08×, p=0.054), but MoE exhibits highly significant overhead: BitNet-Q4/8/16-QMoE-C +11% (p0.001), BitNet-Q4/8-QMoE-C +10% (p0.001). Our most robust finding concerns variance reduction via Levene’s test: curiosity routing achieves 80% F1 variance reduction at () and 94% at () on ESC-50; Quinn shows 50%/85% respectively. Uniform MoE shows higher variance (p=0.851), confirming curiosity-based load balancing essential for predictability.
Effect Sizes and Variability. Narrow confidence intervals (±0.004-0.033 F1) indicate high precision. Non-significance of 4-bit vs 16-bit (0.001 F1 difference, p=0.873 on ESC-50) demonstrates practical equivalence. Cohen’s d ranges 0.42-1.24 for significant comparisons (2-bit: d=-1.24). MoE shows 2-3 higher fold-level variability than single models due to data-dependent routing, but performance rankings remain stable (Spearman’s =0.89, p0.001).
Deployment Implications. Statistical analysis confirms: (1) 4-bit/8-bit achieve practical equivalence with full precision, (2) MoE lacks significant accuracy gains but introduces latency overhead, (3) curiosity routing reduces variance by 50–94% on ESC-50/Quinn (), and (4) adaptive precision allocation () is the most robust benefit, with Q8 automatically receiving uncertain samples. This guides practitioners toward Q4/Q8-Base for most deployments, reserving MoE-Curiosity for safety-critical applications requiring predictable response time and per-sample precision adaptation.
| Test | Comparison | p-value | Sig. |
| F1-Score Improvements | |||
| Cross-Dataset | Q16 vs FP32 (Urban8K) | 0.022 | * |
| Cross-Dataset | Q8/16-PTQ-C vs FP32 (Quinn) | 0.021 | * |
| Ablation | 2-bit vs 16-bit (ESC-50) | 0.004 | ** |
| Ablation | 4-bit vs 16-bit (ESC-50) | 0.873 | ns |
| Latency Overhead | |||
| MoE vs Q8 | Q4/8/16-MoE-C vs Q8-Base | 0.001 | *** |
| MoE vs Q8 | Q4/8-MoE-C vs Q8-Base | 0.001 | *** |
| Speedup | Q1 vs FP32 | 0.054 | ns |
| Variance Reduction (Levene’s test, ESC-50/Quinn) | |||
| Stability | Curiosity (85–94% red.) | 0.001 | *** |
| Stability | Curiosity (50–80% red.) | 0.001 | *** |
| Stability | Q8-Base (68% reduction) | 0.030 | * |
| Confidence-Based Expert Specialization | |||
| Routing | Q8 vs BitNet/Q4 confidence | 0.001 | *** |
6 Discussion
Our findings demonstrate that the primary value of curiosity-driven MoE lies in stability and adaptive precision allocation, not raw accuracy. Simple Q4-Base achieves 0.858 F1, outperforming all MoE configurations on aggregate accuracy. However, curiosity routing provides capabilities unavailable with fixed quantization: 50–94% F1 variance reduction on ESC-50/Quinn () ensures predictable performance, while the Q8 expert automatically receives uncertain samples (20% lower confidence, ), providing per-sample precision adaptation. UrbanSound8K shows no variance benefit (baseline std=0.007), confirming the mechanism targets genuine epistemic uncertainty. For safety-critical applications such as hearing aids or autonomous vehicle perception, this interpretable, uncertainty-aware routing is essential—knowing which expert processed each input and why matters more than marginal accuracy gains.
The counter-intuitive energy pattern (8-bit consuming 2.7 more energy than 16-bit) reflects ARM’s NEON SIMD instruction set, which natively optimizes 16/32-bit operations. 8-bit inference requires additional pack/unpack overhead absent from native precision paths, a platform-specific effect that would differ on hardware with dedicated INT8 accelerators (e.g., edge TPUs, NVIDIA Tensor Cores). Cross-dataset robustness highlights ESC-50’s sensitivity to diverse sounds, while UrbanSound8K’s tolerance supports aggressive quantization for urban monitoring.
Limitations include single-device (Apple M3 CPU) evaluation; multi-platform testing on GPUs, edge TPUs, and other ARM variants could reveal broader gains and different energy profiles. Our evaluation focuses on audio classification using ESC-50 and UrbanSound8K—standard benchmarks appropriate for edge deployment validation. The uncertainty-based routing mechanism (Eq. 8) is domain-agnostic, operating on routing probabilities rather than audio-specific features. Since spectrograms are processed as images, the framework naturally extends to vision tasks; we commit to discussing extensions to other modalities and larger-scale models. Ethically, our low-carbon focus promotes sustainable AI, but deployment in surveillance raises privacy concerns, mitigated by on-device processing.
7 Conclusion
We presented a curiosity-driven quantized MoE framework achieving 50–94% F1 variance reduction on challenging datasets and adaptive precision allocation (). While 4-bit quantization remains optimal for raw accuracy (0.858 F1), curiosity routing provides interpretable, uncertainty-aware expert selection critical for safety-sensitive deployments. Variance reduction scales with baseline uncertainty, confirming the mechanism targets genuine epistemic complexity. Future work includes multi-platform evaluation and extension to vision and multimodal tasks.
Supplementary Material
Appendix A Carbon Emissions Analysis
We provide detailed carbon emission measurements using CodeCarbon Courty et al. (2024) during training and validation on Apple M3 Max CPU (36GB RAM). Table 7 presents emission data for all model configurations.
A.1 Training vs. Deployment Emissions
Single models (1.2M params) emit 22.8 to 55.0 g CO2 during training, while MoE architectures (2.6 to 5.0M params) produce 77.6 to 230.6 g. However, deployment emissions dominate at scale. For 1M daily inferences over 1 year (grid: 367 g CO2/kWh), FP32-Base training emits 22.8 g but deployment produces 2.4 g/year, a 105,000 increase. Q4-Base shows similar patterns: 52.3 g training versus 4.6 g/year deployment. This demonstrates that deployment emissions dominate by approximately 10,000 for models serving more than 1M inferences.
A.2 Green AI Recommendations
Based on our analysis: (1) Use FP32/Q16 for prototyping (23-28 g/experiment); (2) Deploy Q4/Q8 for production to minimize long-term emissions; (3) Schedule training during low-carbon grid hours for 2-5 reduction; (4) Train once and fine-tune for multiple tasks to amortize carbon costs.
| Model | P. | CO2 | Rate | F1 |
| (M) | (g) | (g/s) | ||
| FP32 | 1.19 | 22.85.0 | 0.960.27 | 0.855 |
| Q8-PTQ | 1.19 | 25.917.3 | 1.010.80 | 0.854 |
| Q16 | 1.2 | 27.58.3 | 0.870.35 | 0.859 |
| Q1 | 1.2 | 38.89.6 | 1.060.34 | 0.851 |
| Q2 | 1.2 | 46.414.3 | 1.080.45 | 0.833 |
| BitNet | 1.2 | 48.119.0 | 1.040.57 | 0.793 |
| Q4 | 1.2 | 52.312.8 | 1.050.40 | 0.858 |
| Q8 | 1.2 | 55.016.9 | 1.050.44 | 0.858 |
| Q8PTQ-MoE | 2.59 | 77.637.1 | 1.090.74 | 0.801 |
| Q8/16PTQ-MoE | 4.99 | 83.635.5 | 1.100.64 | 0.820 |
| Q8PTQ-MoE-C | 2.59 | 97.639.8 | 1.040.61 | 0.832 |
| Q8/16PTQ-MoE-C | 4.99 | 105.234.4 | 1.050.51 | 0.832 |
| Q4/8/16-MoE-C | 4.99 | 196.853.1 | 1.060.41 | 0.842 |
| Q4/8-MoE-C | 3.8 | 207.553.2 | 1.110.40 | 0.843 |
| Q4/8/16-MoE | 4.99 | 230.683.5 | 1.090.56 | 0.849 |
Appendix B Model Size Analysis
Table 8 presents model sizes across all quantization schemes. Q8-Base-PTQ achieves highest compression (15.65) through bitwise operations, while Q1-Base reaches 19.99 via BitLinear quantization.
B.1 Optimal Configuration
Q4-Base represents the optimal deployment point with 7.03 compression (687 KB) and 0.858 F1 score, within 0.1% of 16-bit performance while consuming 31% less energy than 8-bit. This balance makes it suitable for most edge deployment scenarios.
B.2 MoE Trade-offs
MoE models show negative compression (0.85) due to multiple expert networks. BitNet-Q4/8/16-QMoE contains 5.67 MB versus single models’ <1.3 MB. However, MoE architectures enable dynamic precision routing, modular updates, and interpretability. For applications prioritizing minimal memory footprint, single quantized models (Q4/Q8-Base) remain superior. MoE is appropriate when interpretability or continual learning are critical.
| Model | Quant. | Bits | Params | KB | MB | Red. |
| FP32 | None | 32 | 1,206,770 | 4,830.8 | 4.83 | 1.00 |
| Q1 | BitLinear | 1 | 1,211,122 | 241.6 | 0.24 | 19.99 |
| Q2 | BitLinear | 2 | 1,211,122 | 390.1 | 0.39 | 12.38 |
| Q4 | BitLinear | 4 | 1,211,122 | 687.1 | 0.69 | 7.03 |
| Q8 | BitLinear | 8 | 1,211,122 | 1,281.0 | 1.28 | 3.77 |
| Q16 | BitLinear | 16 | 1,211,122 | 2,468.8 | 2.47 | 1.96 |
| BitNet | Ternary | 2 | 1,211,122 | 393.9 | 0.39 | 12.26 |
| Q8-PTQ | Bitwise | 2 | 1,205,760 | 308.6 | 0.31 | 15.65 |
| MoE | Hetero. | Mix | 1,415,687 | 5,669.8 | 5.67 | 0.85 |
| MoE-C | H.+Bay. | Mix | 1,415,687 | 5,669.8 | 5.67 | 0.85 |